# Onto-LLM-TAMP: Knowledge-oriented Task and Motion Planning using Large Language Models

### v2

@@ I Introduction
Figure 1: Task execution using the Onto-LLM-TAMP approach. The prompt is processed through the Prompt Tuning module, which uses knowledge-based reasoning to elaborate the user prompt. The state of the environment is textualized using ontological knowledge by the State Descriptor. This detailed elaboration enables the LLM to generate a semantically correct symbolic plan.
@@
To address the above-stated challenges, this study proposes an Onto-LLM-TAMP approach that utilizes knowledge-based reasoning to refine and expand the user’s prompt within the task context. Building on the baseline LLM3 TAMP framework [9], we extend it with our novel Onto-LLM-TAMP method. Some other approaches, such as [13], have also proposed to integrate ontological knowledge with LLM-based task planning, but their primary focus is on object properties to define optimal grasping strategies (e.g., glass $\rightarrow$ pick gently, wood $\rightarrow$ pick stably). On the contrary, we employ knowledge-based reasoning for prompt tuning to refine the action sequence for correct temporal goal ordering, integrating it with the user’s prompt input for the LLM. This ensures that the generated plan is not only correct but also semantically accurate. For example, consider the scenario depicted in Fig. 1, where the user instructs to ”put the banana, apple, and bowl in the plate”. In most cases, the symbolic baseline planner generates a sequence like ”pick banana and apple, and place in the plate” followed by ” pick bowl and place in the plate”. Although it is syntactically correct, it is semantically flawed, as it could result in placing the bowl on top of the apple and banana. Our proposed Onto-LLM-TAMP approach addresses this by incorporating domain-specific knowledge. It integrates a guiding note into the prompt, specifying that ”put bowl before food items because crockery has priority over food items. Put banana and apple after crockery because food items have less priority.” This extended prompt will help the LLM to generate a semantically correct and consistent plan.
@@ IV-A Ontological Knowledge Representation
[…]
@@ V-B Onto-LLM-TAMP Evaluation
During the real-world demonstration, we execute the task of put a banana, an apple, and a bowl in the plate, showcasing the system’s ability to accurately detect, localize, and manipulate objects in a real environment. Figure 8 presents a series of screenshots capturing key moments of the execution process. To ensure precise object identification and spatial positioning, we utilize YOLO-V8 for object detection and FoundationPose for pose estimation. YOLO-V8 enables real-time recognition of objects in the scene, while FoundationPose estimates their positions and orientations, allowing the robot to plan and execute accurate pick-and-place actions. Figures 10 and 9 illustrate the intermediate steps, including detected objects and their estimated poses, demonstrating the effectiveness of our approach in real-world conditions.
@@ VI Conclusions
Acknowledgement
This work has been supported by the Khalifa University of Science and Technology under Award No. RC1-2018-KUCARS, and by the European Commission’s Horizon Europe Framework Programme with the project IntelliMan (AI-Powered Manipulation System for Advanced Robotic Service, Manufacturing and Prosthetics) under Grant Agreement 101070136.

### v3

# LLM-Guided Task and Motion Planning using Knowledge-based Reasoning

@@ I Introduction
To address the above-stated challenges, this study proposes an Onto-LLM-TAMP approach that utilizes knowledge-based reasoning to refine and expand the user’s prompt within the task context. Building on the baseline LLM3 TAMP framework [9], we extend it with our novel Onto-LLM-TAMP method. Some other approaches, such as [13], have also proposed to integrate ontological knowledge with LLM-based task planning, but their primary focus is on object properties to define optimal grasping strategies (e.g., glass $\rightarrow$ pick gently, wood $\rightarrow$ pick stably). On the contrary, we employ knowledge-based reasoning for prompt tuning to refine the action sequence for correct temporal goal ordering, integrating it with the user’s prompt input for the LLM. This ensures that the generated plan is not only correct but also semantically accurate. For example, consider the scenario , where the user instructs to ”put the banana, apple, and bowl in the plate”. In most cases, the symbolic baseline planner generates a sequence like ”pick banana and apple, and place in the plate” followed by ” pick bowl and place in the plate”. Although it is syntactically correct, it is semantically flawed, as it could result in placing the bowl on top of the apple and banana. Our proposed Onto-LLM-TAMP approach addresses this by incorporating domain-specific knowledge. It integrates a guiding note into the prompt, specifying that ”put bowl before food items because crockery has priority over food items. Put banana and apple after crockery because food items have less priority.” This extended prompt will help the LLM to generate a semantically correct and consistent plan.
@@ IV-A Ontological Knowledge Representation
$\text{{\color[rgb]{0.64,0.08,0.98}\definecolor[named]{pgfstrokecolor}{rgb}{0.64,0.08,0.98}Rule1}}\sqsubseteq\text{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}ActionPriority}}\sqcap$ $\exists\text{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}hasAction}.{\color[rgb]{0.64,0.08,0.98}\definecolor[named]{pgfstrokecolor}{rgb}{0.64,0.08,0.98}Put}}$
$\sqcap\exists\text{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}hasObjectType}.{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}Crockery}}\sqcap$ $\exists\text{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}hasPriority}}.1$ (3)
$\text{{\color[rgb]{0.64,0.08,0.98}\definecolor[named]{pgfstrokecolor}{rgb}{0.64,0.08,0.98}Rule2}}\sqsubseteq\text{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}ActionPriority}}\sqcap$ $\exists\text{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}hasAction}.{\color[rgb]{0.64,0.08,0.98}\definecolor[named]{pgfstrokecolor}{rgb}{0.64,0.08,0.98}Put}}$
$\sqcap\exists\text{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}hasObjectType}.{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}FoodItem}}\sqcap$ $\exists\text{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}hasPriority}}.2$ (4)
@@ V-B Onto-LLM-TAMP Evaluation
During the real-world demonstration, we execute the task of put a banana, an apple, and a bowl in the plate, showcasing the system’s ability to accurately detect, localize, and manipulate objects in a real environment. Figure 7 presents a series of screenshots capturing key moments of the execution process. To ensure precise object identification and spatial positioning, we utilize YOLO-V8 for object detection and FoundationPose for pose estimation. YOLO-V8 enables real-time recognition of objects in the scene, while FoundationPose estimates their positions and orientations, allowing the robot to plan and execute accurate pick-and-place actions. Figures 9 and 8 illustrate the intermediate steps, including detected objects and their estimated poses, demonstrating the effectiveness of our approach in real-world conditions.